Astronomy Report – March 2nd

Crew 223 Astronomy Report Mar 02 2020

Name: Florian Delpech
Crew: 223
Date: 3/2/2020

MDRS ROBOTIC OBSERVATORY
Robotic Telescope Requested (choose one) MDRS-WF

Objects to be Imaged this Evening: M81, M106, NGC2403, NGC6946, IC1805. This is part of the astrophotography project

Images submitted with this report: N/A

Problems Encountered: N/A

MUSK OBSERVATORY

Solar Features Observed: I used for the first time the observatory today following the quick guide. I tried to observe visually some solar features but I did not notice any special phenomenon.

Images submitted with this report: N/A

Problems Encountered: N/A

Operations Report – March 1st

Crew 223 Operations Report 03/01/2020

SOL: 0

Name of person filing report: Luc FORTIN

Non-nominal systems: Nothing to report

Notes on non-nominal systems: Nothing to report

Saturday generator:
Off at 5:30am
On at 8:00pm SOC was 80%

Sunday generator:
Off at 5:30am
On at 7:00pm SOC was 79%

Diesel Reading – 70%

Station Propane Reading – 82%

Ethanol Free Gasoline – 7 gallons

Water (loft tank): 47 gallons

Water Meter: 149647.32 units

Water (static tank) – 381 gallons

Static to Loft Pump used – yes

Water in GreenHab – 235.2 gallons

Water in ScienceDome: 0 gallons

Toilet tank emptied: no

Deimos rover used: yes (for training with Shannon)
Hours: 147
Beginning charge: 100%
Ending charge: Non checked
Currently charging: yes

Sojourner rover used: ASSIGNED TO DIRECTOR
Hours: Nothing to report
Beginning charge: Nothing to report
Ending charge: Nothing to report
Currently charging: Nothing to report

Spirit rover: yes (for training with Shannon)
Hours: 133.2
Beginning charge: 100%
Ending charge: Non checked
Currently charging: yes

Opportunity rover used: moved from campus to the shop to repair the brakes. Hours: Nothing to report
Beginning charge: Nothing to report
Ending charge: Nothing to report
Currently charging: Nothing to report

Curiosity rover used: yes (for training with Shannon)
Hours: 139.1
Beginning charge: 100%
Ending charge: Non checked
Currently charging: yes

Notes on rovers: Nothing to report

ATV’s Used: yes (Honda, 1, 2 and 3)
Reason for use: Training with Shannon
Oil Added? no
ATV Fuel Used: almost 2% of the tank.
Hours the ATVs were Used today: 5 minutes
Notes on ATVs: As it already happened with the previous crew, the Honda ATV was stuck in 5th gear and refused to go back into neutral (reason = slowing down while 5th gear engaged). Shannon fixed it by moving manually the ATV. The gears are now working perfectly.

HabCar used and why, where? We used the Habcar to go to Hanksville to fill up the GreenHab tank (which was full at noon). We also refueled it. Habcar gas tank cap broken. The HabCar is now parked at the MDRS.

CrewCar used and why, where? The CrewCar is in Grand Junction since this morning, thanks to crew 222.

General notes and comments: Nothing to report

Summary of internet: All ok.

Summary of suits and radios: Suit 3 was used during the engineer EVA.

Summary of Hab operations: Nothing to report

Summary of GreenHab operations: Nothing to report

Summary of ScienceDome operations: Nothing to report

Summary of RAM operations: Nothing to report

Summary of any observatory issues: Nothing to report

Summary of health and safety issues: Nothing to report

Questions, concerns and requests to Mission Support: Nothing to report

Operations Report – March 2nd

Crew 223 Operations Report 03/02/2020

SOL: 1

Name of person filing report: Luc FORTIN

Non-nominal systems: Nothing to report

Notes on non-nominal systems: Nothing to report

Sunday generator:
Off at 5:30am
On at 7:00pm SOC was 79%

Monday generator:
Off at 4:30am
On at 9pm SOC was XX%

Diesel Reading – 70%

Station Propane Reading – 80%

Ethanol Free Gasoline – 7 gallons

Water (loft tank): 47 gallons

Water Meter: 149665.40 units

Water (static tank) – 363 gallons

Static to Loft Pump used – yes

Water in GreenHab – 218.3 gallons

Water in ScienceDome: 0 gallons

Toilet tank emptied: yes

Deimos rover used: yes
Hours: 149.1
Beginning charge: 100%
Ending charge: 100%
Currently charging: yes

Sojourner rover used: ASSIGNED TO DIRECTOR
Hours: Nothing to report
Beginning charge: Nothing to report
Ending charge: Nothing to report
Currently charging: Nothing to report

Spirit rover: no
Hours: 133.2
Beginning charge: Nothing to report
Ending charge: Nothing to report
Currently charging: yes

Opportunity rover used: moved from campus to the shop to repair the brakes. Hours: Nothing to report
Beginning charge: Nothing to report
Ending charge: Nothing to report
Currently charging: Nothing to report

Curiosity rover used: yes
Hours: 139.2
Beginning charge: 100%
Ending charge: 100%
Currently charging: yes

Notes on rovers: Nothing to report

ATV’s Used: yes (Honda, 1, 2 and 3)
Reason for use: Run during the engineer EVA for battery
Oil Added? no
ATV Fuel Used: Nothing to report
Hours the ATVs were Used today: 5 minutes
Notes on ATVs: The ATV 1’s oil light up.

HabCar used and why, where? Nothing to report.

CrewCar used and why, where? The CrewCar is in Grand Junction.

General notes and comments: Nothing to report

Summary of internet: All ok.

Summary of suits and radios: Suits 4, 7, 8, 9 were used during the EVA. We had troubles with suit number 7 (intermittent fan, same as crew 222 and 221), I tried to fix it but didn’t manage to. I just find an email sent last friday by Scott Davis giving some ideas to fix it, so I will try again tomorrow.

Summary of Hab operations: Nothing to report

Summary of GreenHab operations: Nothing to report

Summary of ScienceDome operations: Nothing to report

Summary of RAM operations: Tried to fix suit 7 without success.

Summary of any observatory issues: Nothing to report

Summary of health and safety issues: Nothing to report

Questions, concerns and requests to Mission Support: Nothing to report

Operations Report – March 2nd

Crew 223 Operations Report 03/01/2020

SOL: 0

Name of person filing report: Luc FORTIN

Non-nominal systems: Nothing to report

Notes on non-nominal systems: Nothing to report

Saturday generator:
Off at 5:30am
On at 8:00pm SOC was 80%

Sunday generator:
Off at 5:30am
On at 7:00pm SOC was 79%

Diesel Reading – 70%

Station Propane Reading – 82%

Ethanol Free Gasoline – 7 gallons

Water (loft tank): 47 gallons

Water Meter: 149647.32 units

Water (static tank) – 381 gallons

Static to Loft Pump used – yes

Water in GreenHab – 235.2 gallons

Water in ScienceDome: 0 gallons

Toilet tank emptied: no

Deimos rover used: yes (for training with Shannon)
Hours: 147
Beginning charge: 100%
Ending charge: Non checked
Currently charging: yes

Sojourner rover used: ASSIGNED TO DIRECTOR
Hours: Nothing to report
Beginning charge: Nothing to report
Ending charge: Nothing to report
Currently charging: Nothing to report

Spirit rover: yes (for training with Shannon)
Hours: 133.2
Beginning charge: 100%
Ending charge: Non checked
Currently charging: yes

Opportunity rover used: moved from campus to the shop to repair the brakes. Hours: Nothing to report
Beginning charge: Nothing to report
Ending charge: Nothing to report
Currently charging: Nothing to report

Curiosity rover used: yes (for training with Shannon)
Hours: 139.1
Beginning charge: 100%
Ending charge: Non checked
Currently charging: yes

Notes on rovers: Nothing to report

ATV’s Used: yes (Honda, 1, 2 and 3)
Reason for use: Training with Shannon
Oil Added? no
ATV Fuel Used: almost 2% of the tank.
Hours the ATVs were Used today: 5 minutes
Notes on ATVs: As it already happened with the previous crew, the Honda ATV was stuck in 5th gear and refused to go back into neutral (reason = slowing down while 5th gear engaged). Shannon fixed it by moving manually the ATV. The gears are now working perfectly.

HabCar used and why, where? We used the Habcar to go to Hanksville to fill up the GreenHab tank (which was full at noon). We also refueled it. Habcar gas tank cap broken. The HabCar is now parked at the MDRS.

CrewCar used and why, where? The CrewCar is in Grand Junction since this morning, thanks to crew 222.

General notes and comments: Nothing to report

Summary of internet: All ok.

Summary of suits and radios: Suit 3 was used during the engineer EVA.

Summary of Hab operations: Nothing to report

Summary of GreenHab operations: Nothing to report

Summary of ScienceDome operations: Nothing to report

Summary of RAM operations: Nothing to report

Summary of any observatory issues: Nothing to report

Summary of health and safety issues: Nothing to report

Questions, concerns and requests to Mission Support: Nothing to report

Greenhab Report – March 2nd

Crew 223 GreenHab Report – 02-03-2020 GreenHab Officer: Valentin BELLEMIN-LAPONNAZ

Environmental control: Ambient with Door open / Door closed

Average temperature: 23.2°C
Low temperature: 19°C
High temperature: 30°C

Hours of supplemental light: 4 hours (from 9P.M to 1 a.m.)

Daily water usage for crops: 16.8 gallons

Daily water usage for research and/or other purposes: 0.5 gallon from loft tank

Water in Blue Tank— 218.3 gallons

Time(s) of watering for crops: 3 (one only for tomatoes)

Changes to crops: I planted lettuces in the pots where lettuces had been harvested yesterday. I transplanted zucchini plants that ware in common pots into two separated pots to allow them to grow better.

Narrative: This morning, I clarified tomatoes to give them more space. I also planted lettuces, as I harvested some yesterday. For lunch time, I harvested zucchini flower to make an Italian recipe and basil to make homemade fresh green pesto. This afternoon, we took care of spirulina with our commander. We refilled the left tank and fed algae them with nutrients. After this, I worked on my personal experiment : Music For Plants. I moisturized dry watercress seeds with Murashige and Skoog to plant them tomorrow. After watering the plants at the end of the afternoon, I harvested Zebra Mint to do some moroccan tea.

Harvest: 6g Zucchini Flower / 18g Basil / Spirulina / 8g Zebra Mint

Support/supplies needed: NA

Journalist Report – March 2nd

Crew 223 Journalist Report 02Mar2020
Author: Clément Plagne, Crew Journalist
Sol 1
Title: Baby steps

After years spent on Earth, with the comfort of modern technology and the plentiful Western lifestyle, seeing ourselves so limited is as refreshing as it is a shock. We’re like born again on Mars, having to re-learn everything as we move along in our new home. And, as newborns, even our first steps weren’t so easy. Today was EVA day, and it was time for four of us to step out into the (nearly) unknown, making our way into the hills after a first run to replace the batteries of our atmospheric experiments. Between the heat outside, the weight of the life support system and the little air we got from it, it was no walk in the park. Still, with every heavy step, we discovered beautiful features of the terrain, and it was like facing a brand-new landscape every few meters. Deserts are not boring, that’s for sure.

Like the explorers of old, we were on our own, too far away to communicate with help of any kind, and having to figure things out ourselves if anything went wrong. During the EVA, Blandine’s suit had issues with its ventilation systems – while she was perfectly safe in earth’s atmosphere, this could have proven problematic or even critical in real conditions. We had to react quickly, and solved the issue even with our unwieldy gloves. It’ll be up to Luc, the Crew Engineer, to figure out what is wrong and solve the issues.

While we were out, something possibly even more important was going on: Valentin, our botanist and resident cook for the day, was starting to get busy in the kitchen. For lunch, he was preparing homemade pasta with pesto, made from basil harvested this very morning in the GreenHab. As a side, zucchini flowers dipped in batter and shallow fried. An absolute success! And, as if that wasn’t enough, bread was rising for tonight’s dinner. Food: truly the best morale booster for an all-French crew!

Scarcity breeds creativity, so here’s to many more days of delicious food and creative problem solving!

EVA Report – March 2nd

Crew 223 EVA Report 02-03-2020
EVA # 2
Author: Aurélien Mure (Commander) and Blandine Gorce (XO)
Purpose of EVA: Change batteries of LOAC and MegaAres, and then first exploration for the crew 223
Start time: 9:14 AM
End time: 10:57 AM
Narrative: First, we went out to check all the rovers and the water in teams of two. We took then the two rovers to go on our experiments site to change
batteries. Everything went fine so we moved a few meters to go to our site of exploration: the Marble’s Ritual. We arrived there at 09:43 to start our exploration, Curiosity was at 96% of charge, and Deimos was at 100%.
All of us were very excited for that. We walked for approximately one hour between Marble’s Ritual and Phobo’s Peak. Landscapes were really beautiful and we were amazed to see so many different types and rocks. We walked around to appreciate the landscape. Then we came back to the hab around 10:45. We entered the airlock again at 10:52 AM for our 5 minutes pressurization.
Destination: Marble’s Ritual
Coordinates (use UTM NAD27 CONUS):4250800/518500, 4251750/518700 Participants: Team red: Aurélien Mure (Commander), Clément Plagne (Journalist), Team Blue: Luc Fortin (Engineer), Blandine Gorce (HSO) Road(s) and routes per MDRS Map: Entrance Road, Cow Dung Road Mode of travel: Rovers : Curiosity & Deimos

Sol Summary – March 2nd

Crew 223 Sol Summary Report 02-03-2020

Sol: 1 Summary Title: Giant leap into martian routine

Author’s name: Aurélien Mure – Crew Commander

Mission Status: Crew 223 discovering sim

Sol Activity Summary:

The crew woke up at 6:50am as it will during the whole mission. The sport session began at 7:00am. It was the first long session for 5 members of the crew. They were really motivated and did a great workout session. We went through the breakfast as a large family! The crew took time to prepare the first EVA for an hour. In fact, Supaero crews have special EVA checklist and protocols taking time. Blandine, Clément, Luc and I went out to change the LOAC and MegaAres batteries and start exploring. We went to Marble’s Ritual and explored until Phobos Peak. It was a great training EVA and the whole crew was perfectly handling the operations.
For lunch Valentin cooked homemade pasta with fresh pesto sauce and deep fried zucchini flowers.
I think we are going to spend two amazing weeks in terms of food! In the afternoon each crew member worked on its own experiments or reports after having spent half an hour playing cards together. Then the crew started the language human factor experiment at 4:00pm. We went through the training to emergency equipment Blandine gave us.

Look Ahead Plan:

Tomorrow we will go on EVA in the morning if the weather allows it.

The EVA will be focused on familiarizing the rest of the crew to the protocoles and to the routine of battery replacement for the LOAC and MegaAres.

Anomalies in work: None

Weather: Sunny and warm all day (-2°C at 7:00am, about 11°C at 2:00pm)

Crew Physical Status: All crew members safe and sound

EVA: NA

Reports to be file: Journalist Report, Operations report, Food Inventory, EVA Request, Green Hab report

Support Requested: None